OpenStudy (anonymous):

slope intercept form for 3x - 5y =25

OpenStudy (petiteme):

slope intercept form is y=mx+b so 3x - 5y =25 transpose 3x to the other side -5y= 25 - 3x divide both sides by -5 so y= -5 + 3x/5

OpenStudy (anonymous):

thanks so much!

OpenStudy (petiteme):

no prob :) you're welcome! :)